# Lightweight posterior construction for gravitational-wave catalogs with the Kolmogorov-Arnold network

### v1

@@ III Case Tests
Before applying the KAN-based neural density estimator to real GW posterior data, we construct three illustrative cases to test the reliability of our model and explore its ability to fit posterior distributions into analytic expressions. We choose a unimodal distribution, a conditional unimodal distribution, and a bimodal distribution as our test cases Papamakarios et al. (2017). In each example, we start with the ground-truth PDF to obtain the raw samples, and then use these as input for the KAN-based neural density estimator to fit for the distribution’s analytic form. It is important to note that we do not specify the analytic form of the distribution in advance. Instead, we allow the KAN-based neural density estimator to freely explore the fitting expressions without any additional input information from human.
@@
Its distribution forms a crescent shape, and we aim to use this example to test the method’s capability for non-Gaussian distributions. The PDF of the second example, a conditional unimodal distribution, reads
@@
This example involves a highly distorted bimodal distribution, which is used to test the neural density estimator’s fitting ability with multimodal distributions.
@@ IV Lightweight posterior construction for GW catalogs
In this section, we apply the KAN-based neural density estimator to real GW data. We select six GW parameters—chirp ${\cal M}, mass$ ratio $q$, two spin magnitudes $\chi_{1}$ and $\chi_{2}$, and two spin inclinations $\theta_{1}$ and $\theta_{2}$—and attempt to encode the posterior distributions of ten GW events: GW150914, GW151012, GW151226, GW170104, GW170608, GW170729, GW170809, GW170814, GW170818, and GW170823 Abbott et al. (2024). Brief information of these events is presented in Appendix A. The posterior samples are obtained from the Gravitational Wave Open Science Center (GWOSC) [^1] https://gwosc.org/ Abbott et al. (2021e, 2023d); Abac et al. (2025d). We evaluate the reliability of our two encoded data products—high-fidelity neural network weights and analytic expressions—by comparing the posteriors of raw samples with those of the resampled samples. Additionally, we further ensure the reliability of encoded data products in practical applications by performing population inference using reconstructed posterior samples.
@@
where $\mu[\ln{\mathcal{M}}]$ and $\sigma[\ln{\mathcal{M}}]$ are the mean and standard deviation of the logarithm of the chirp mass respectively. Among the remaining parameters, the mass ratio and spin magnitudes are constrained within the interval $[0,1]$, while the spin inclinations are restricted to $0,\pi]$. We follow Hoy and Raymond (2021) and apply the logit transformation to these parameters,
@@
Here, $\theta$ represents the original parameter value, and $a$ and $b$ are the lower and upper bounds of the parameter, respectively. All these transformations are invertible and performed before training the neural network. After the training phase, inverse transformations are performed during the sampling phase to recover the original parameter values. Other methods such as boundary reflection can also be used to handle the bounded parameters Jones (1993) .
@@
After data preprocessing, neural network training, and symbolification, we obtain two encoded representations of the posterior samples for each event. In our model for these ten GW events, all neural networks adopt a two-layer architecture, with five Gaussian components used to fit each conditional probability distribution, and each unmasked edge contains twenty-three cubic B-spline functions. Taking GW150914 as an example, Table 1 lists the storage requirements and sampling rates for both types of encoded data products. The neural network weights and analytic expressions achieve approximately $90\times$ and $3500\times$ compression ratios in storage, respectively.
@@
The resampling results are shown in Fig. 4 and Fig. 5. As shown in Fig. 4, the samples generated from the neural network weights exhibit close agreement with the original distribution. In Fig. 5, samples generated with the analytic expressions capture the overall structure of the posterior distribution, but exhibit some deviations from the original distributions. Compared with the neural network weights, this reflects some trade-off between fidelity and the gains in storage efficiency, interpretability, and expression simplicity. By replacing the network’s nonlinear edge functions with more restrictive analytical forms, the model’s complexity is inevitably reduced and some fidelity in capturing local features is lost. The analytic expression data products occupy only $\mathcal{O}\,(1\,{\rm KB})$ of storage. In Fig. 5 we also show samples from a fitted multivariate Gaussian distribution, where we use the same data preprocessing method as in the KAN-based neural density estimator. In preparing them, we start from the initial mean and covariance matrix of the preprocessed raw samples, and obtain the optimal mean and covariance matrix of the 6-dimensional Gaussian distribution by minimizing the average negative log-posterior. Compared to other representations, such as summarizing GW posteriors by their mean and covariance matrix, as well as its higher-order extension Wang et al. (2022), the analytic expressions produced by the method of this work are much better at capturing the distortions and multimodal features of GW posterior distributions, giving a better balance between accuracy and storage efficiency. The probability-probability (P-P) plot comparing the raw and resampled samples of GW150914 is shown in Fig. 6. The resampling results for other nine GW events are shown in Appendix. A.
@@
To quantify the differences between the distributions of the raw and resampled GW150914 posterior samples, we calculate the Wasserstein distances Vaserstein (1969); Peyré and Cuturi (2019) of the six one-dimensional marginalized distributions between raw and reproduced posteriors. Wasserstein distance is a distance function defined between two probability distributions that measure the similarity of them. Compared with the more commonly used Kullback-Leibler divergence Kullback and Leibler (1951), Wasserstein distance is more universal and can reflect the similarity between two distributions which overlap very little. Intuitively, if one regards each probability distribution as an equal (normalized) pile of soil, then the Wasserstein distance can be interpreted as the required minimum “work” to deform one pile of soil into another, which is equal to the amount of soil to be moved multiplied by the average “distance” between two piles (distributions) Wang et al. (2022). We normalize the Wasserstein distances using the standard deviation of each one-dimensional marginalized distribution in raw posteriors and the results are shown in Tab. 2. Similar distributions will have a small Wasserstein distance. We can see that most of the one-dimensional marginalized distributions between raw samples and samples resampled with our data products match very well with Wasserstein distances smaller than $0.02$, indicating that the difference is less than $0.02$ standard deviation. Although samples generated with analytic expressions show relatively larger deviation in the marginalized distribution of $a_{2}$, analytic expressions outperform the multivariate Gaussian model in all six parameters.
@@
Table 1: Comparison between two data-compression products for GW150914’s posterior distribution.GG” denotes the number of Gaussian components in GMMs that fit the conditional probability distribution (1); “Hidden” indicates the number of hidden layers in the neural network architecture before symbolification; “Storage” shows the storage space ratio of data products to raw samples, with the raw data containing 147,634 six-dimensional posterior samples in HDF5 format; “Sampling Time” represents the average re-sampling duration of 20 runs with 2,000 samples per run.
@@
Many key scientific results in GW astronomy, such as constraints on the mass and spin distributions of black holes Dong et al. (2024); Tong et al. (2022), rely on population inference that combines posterior samples from many events. In the era of next-generation GW detectors, the number of detected events is expected to increase explosively, and the demand for posterior samples in population studies will grow accordingly. This will pose challenges on users for both data storage and transmission. The primary motivation for proposing a lightweight posterior catalog is to address the need for efficient storage and distribution of large-scale GW posterior data in the future. To ensure that our method preserves not only event-level fidelity but also the statistical properties required for population-level scientific applications, we further validate the reliability of our lightweight data products by performing population inference using samples resampled using our two types of compact data products.
@@
where $P(\mathcal{D}|\bm{H})$ is the total marginalized likelihood over the source parameters of the events and $P(\bm{H})$ is the prior distribution of the population parameters. Using posterior samples in the GW catalog, $P(\mathcal{D}|\bm{H})$ is computed as follows Thrane and Talbot (2019),
@@
Here, the summation is over the $n_{i}$ posterior samples, $\{\bm{\theta}_{i}^{k}\}$, for event $i$. The term $\pi(\bm{\theta}_{i}^{k}|\emptyset)$ is the original prior used to generate these samples, while $\pi(\bm{\theta}_{i}^{k}|\bm{H})$ represents the re-evaluated prior for those same samples in the population model.
We compare the posterior distributions, $P(\bm{H}|\mathcal{D})$, based on three sets of samples: the raw samples for each event from GWOSC, the recovered samples using corresponding neural network weights and the recovered samples using corresponding analytic expressions. We use the GWPopulation package Talbot et al. (2025) to perform population inference on these three sets of samples. The analyses use the two-component-primary-mass-ratio and iid-spin models, with detailed description and prior choices of the hyperparameters given in Appendix B. Comparative results for the three population-level posterior distributions are shown in Fig. 7. As shown in the figure, population inference results of raw samples and samples recovered using neural network weights match very well and the result of samples recovered using analytic expressions exhibit some slight deviations, validating the feasibility of our compression strategy.
Figure 7: Marginalized one- and two-dimensional distributions, comparing the population inference results using raw samples (blue), samples generated with our neural networks (orange) and analytic expressions (pink) of 10 GW events. Contour lines in the two-dimensional joint distributions delineate the $50\%$ and $90\%$ credible regions. The sampling is accomplished using the dynesty sampler Speagle (2020) and nlive is set to 2,000.
@@ Acknowledgements
This work was supported by the Beijing Natural Science Foundation (1242018), the National Natural Science Foundation of China (123B2043), the National SKA Program of China (2020SKA0120300), the Max Planck Partner Group Program funded by the Max Planck Society, and the High-performance Computing Platform of Peking University.
@@
Appendix A GW Events and Sampling Results
In this appendix, we give some information of binary black hole GW events in GWTC-1 (GW150914, GW151012, GW151226, GW170104, GW170608, GW170729, GW170809, GW170814, GW170818, GW170823 Abbott et al. (2019a)), and the resampling results obtained with our compact data products.
All ten GW events used in this paper are binary black hole mergers composed of stellar-mass black holes, originally detected during the LIGO/Virgo’s first and second observing runs Abbott et al. (2019a). The raw posterior samples used in this paper are taken from the GWTC-2.1 catalog Abbott et al. (2024), which features a reanalysis of these events with state-of-the-art waveform models. This population covers a range of total masses from approximately $18.5\,M_{\odot}$ (GW170608) to $84.4\,M_{\odot}$ (GW170729). While most systems have comparable component masses, the spin characteristics show important variations. The majority of events have effective inspiral spins ($\chi_{\rm eff}$) consistent with zero. However, GW151226 stands out with a measured positive spin, providing evidence for at least one spinning black hole in that system. Furthermore, parameter uncertainties vary significantly across the events, often correlating with detection significance: the inaugural loud event GW150914 with a network SNR of $26$ has well-constrained properties, whereas the most distant event GW170729 at redshift $z\approx 0.44$ exhibits much broader posterior distributions due to its lower SNR.
The sampling results are shown in Fig. 8 and Fig.9. As shown in Fig. 8, the samples generated from the neural network weights exhibit close agreement with the original distribution from GWOSC. In Fig. 9, samples generated with the analytic expressions capture the overall structure of the posterior distribution, but exhibit some deviations from the original distributions. In producing the results in the figures, the data preprocessing method, neural network configurations, training processes, and training devices of all ten events are kept consistent. During resampling, the number of samples for each event is the same as that of the raw samples, and the resampling device also remains consistent.

### v2

@@ III Case Tests
Before applying the KAN-based neural density estimator to real GW posterior data, we construct three illustrative cases to test the reliability of our model and explore its ability to fit posterior distributions into analytic expressions. We choose a unimodal distribution, a conditional unimodal distribution, and a bimodal distribution as our test cases Papamakarios:2017tec. In each example, we start with the ground-truth PDF to obtain the raw samples, and then use these as input for the KAN-based neural density estimator to fit for the distribution’s analytic form. It is important to note that we do not specify the analytic form of the distribution in advance. Instead, we allow the KAN-based neural density estimator to freely explore the fitting expressions without any additional input information from human.
@@
Its distribution forms a crescent shape, and we aim to use this example to test the method’s capability of accommodating non-Gaussian distributions. The PDF of the second example, a conditional unimodal distribution, reads
@@
This example involves a highly distorted bimodal distribution, which is used to test the neural density estimator’s fitting ability with multimodal distributions. Beyond these examples, we also present one additional illustrative example based on an exponential distribution in Appendix A to more comprehensively demonstrate the effectiveness of this method in modeling diverse distributions.
@@ IV Lightweight posterior construction for GW catalogs
In this section, we apply the KAN-based neural density estimator to real GW data. We select six GW parameters—chirp ${\cal M}, mass$ ratio $q$, two spin magnitudes $\chi_{1}$ and $\chi_{2}$, and two spin inclinations $\theta_{1}$ and $\theta_{2}$—and attempt to encode the posterior distributions of ten GW events: GW150914, GW151012, GW151226, GW170104, GW170608, GW170729, GW170809, GW170814, GW170818, and GW170823 LIGOScientific:2021usb. Brief information of these events is presented in Appendix B. The posterior samples are obtained from the Gravitational Wave Open Science Center (GWOSC) [^1] https://gwosc.org/ LIGOScientific:2019lzm; KAGRA:2023pio; LIGOScientific:2025snk. We evaluate the reliability of our two encoded data products—high-fidelity neural network weights and analytic expressions—by comparing the posteriors of raw samples with those of the resampled samples. Additionally, we further ensure the reliability of encoded data products in practical applications by performing population inference using reconstructed posterior samples.
@@
where $\mu[\ln{\mathcal{M}}]$ and $\sigma[\ln{\mathcal{M}}]$ are the mean and standard deviation of the logarithm of the chirp mass respectively. Among the remaining parameters, the mass ratio and spin magnitudes are constrained within the interval $[0,1]$, while the spin inclinations are restricted to $0,\pi]$. We follow Hoy:2020vys and apply the logit transformation to these parameters,
@@
Here, $\theta$ represents the original parameter value, and $a$ and $b$ are the lower and upper bounds of the parameter, respectively. All these transformations are invertible and performed before training the neural network. After the training phase, inverse transformations are performed during the sampling phase to recover the original parameter values. Other methods such as boundary reflection can also be used to handle the bounded parameters jones1993simple.
@@
After data preprocessing, neural network training, and symbolification, we obtain two encoded representations of the posterior samples for each event. For these ten GW events, all neural networks adopt a compact two-layer architecture. We empirically explored different choices for the number of Gaussian components $G$ and B-spline basis functions $B$. Increasing $G$ and $B$ beyond moderate values yielded only marginal gains, whereas reducing them degraded reconstruction fidelity. Accordingly, we selected $G=5$ and $B=23$ as a practical compromise between accuracy and compactness. Taking GW150914 as an example, Tab. 1 lists the storage requirements and sampling rates for both types of encoded data products. The neural network weights and analytic expressions achieve approximately $90\times$ and $3500\times$ compression ratios in storage, respectively.
@@
The resampling results are shown in Fig. 4 and Fig. 5. As shown in Fig. 4, the samples generated from the neural network weights exhibit close agreement with the original distribution. In Fig. 5, samples generated with the analytic expressions capture the overall structure of the posterior distribution, but exhibit some deviations from the original distributions. Compared with the neural network weights, this reflects some trade-off between fidelity and the gains in storage efficiency, interpretability, and expression simplicity. By replacing the network’s nonlinear edge functions with more restrictive analytical forms, the model’s complexity is inevitably reduced and some fidelity in capturing local features is lost. The analytic expression data products occupy only $\mathcal{O}\,(1\,{\rm KB})$ of storage. In Fig. 5 we also show samples from a fitted multivariate Gaussian distribution, where we use the same data preprocessing method as in the KAN-based neural density estimator. In preparing them, we start from the initial mean and covariance matrix of the preprocessed raw samples, and obtain the optimal mean and covariance matrix of the 6-dimensional Gaussian distribution by minimizing the average negative log-posterior. Compared to other representations, such as summarizing GW posteriors by their mean and covariance matrix, as well as its higher-order extension Wang:2022kia, the analytic expressions produced by the method of this work are much better at capturing the distortions and multimodal features of GW posterior distributions, giving a better balance between accuracy and storage efficiency. The probability-probability (P-P) plot comparing the raw and resampled samples of GW150914 is shown in Fig. 6. The resampling results for other nine GW events are shown in Appendix. B.
@@
To quantify the differences between the distributions of the raw and resampled GW150914 posterior samples, we calculate the Wasserstein distances 1969Markov; computational2019Peyre of the six one-dimensional marginalized distributions between raw and reproduced posteriors. Wasserstein distance is a distance function defined between two probability distributions that measure the similarity of them. It is worth noting that this distance metric characterizes differences in one-dimensional marginal distributions, while corner plots provide a complementary way to compare correlations among parameters. Compared with the more commonly used Kullback-Leibler divergence Kullback:1951zyt, Wasserstein distance is more universal and can reflect the similarity between two distributions which overlap very little. Intuitively, if one regards each probability distribution as an equal (normalized) pile of soil, then the Wasserstein distance can be interpreted as the required minimum “work” to deform one pile of soil into another, which is equal to the amount of soil to be moved multiplied by the average “distance” between two piles (distributions) Wang:2022kia. We normalize the Wasserstein distances using the standard deviation of each one-dimensional marginalized distribution in raw posteriors and the results are shown in Tab. 2. Similar distributions will have a small Wasserstein distance. We can see that most of the one-dimensional marginalized distributions between raw samples and samples resampled with our data products match very well with Wasserstein distances smaller than $0.02$, indicating that the difference is less than $0.02$ standard deviation. Although samples generated with analytic expressions show relatively larger deviation in the marginalized distribution of $\theta_{2}$, analytic expressions outperform the multivariate Gaussian model in all six parameters. As a metric for quantifying differences between high-dimensional distributions, we compute the two-dimensional Hellinger distances hellinger1909neue for pairwise combinations of the parameters of GW150914. The results, summarized in Fig. 7, provide a complementary assessment of the joint correlations.
Additionally, we apply the above KAN-based density estimator to the reconstruction and compression of the full 15-dimensional posterior for GW150914. The data product based on network weights achieves a $34\times$ compression, demonstrating that the method remains robust and effective in the full-parameter scenario. Details of the experimental configuration and the reconstructed posteriors are presented in Appendix C.
@@
Table 1: Comparison between two data-compression products for GW150914’s posterior distribution.GG” denotes the number of Gaussian components in GMMs that fit the conditional probability distribution (1); “Hidden” indicates the number of hidden layers in the neural network architecture before symbolification; “Storage” shows the storage space ratio of data products to raw samples, with the raw data containing 147,634 six-dimensional posterior samples in HDF5 format; “Sampling Time” represents the average re-sampling duration of 20 runs with 2,000 samples per run. Averaged over five runs, network training took 27.14 seconds on the same hardware as in resampling.
@@
Figure 7: Two-dimensional Hellinger distances for pairwise combinations of the parameters of GW150914. The distances take values in $[0,1]$, with smaller values indicating closer agreement. The top right corner shows the comparison between original samples and samples generated using network weights. The bottom left corner shows the comparison between original samples and samples generated using analytic expressions.
Many key scientific results in GW astronomy, such as constraints on the mass and spin distributions of black holes Dong:2024zzl; Tong:2022iws, rely on population inference that combines posterior samples from many events. In the era of next-generation GW detectors, the number of detected events is expected to increase explosively, and the demand for posterior samples in population studies will grow accordingly. This will pose challenges on users for both data storage and transmission. The primary motivation for proposing a lightweight posterior catalog is to address the need for efficient storage and distribution of large-scale GW posterior data in the future. To ensure that our method preserves not only event-level fidelity but also the statistical properties required for population-level scientific applications, we further validate the reliability of our lightweight data products by performing population inference using samples resampled using our two types of compact data products.
@@
where $P(\mathcal{D}|\bm{H})$ is the total marginalized likelihood over the source parameters of the events and $P(\bm{H})$ is the prior distribution of the population parameters. Using posterior samples in the GW catalog, $P(\mathcal{D}|\bm{H})$ is computed as follows thrane2019introduction,
@@
Here, the summation is over the $n_{i}$ posterior samples, $\{\bm{\theta}_{i}^{k}\}$, for event $i$. The term $\pi(\bm{\theta}_{i}^{k}|\emptyset)$ denotes some default prior in the original analysis and $\mathcal{Z}$ denotes the corresponding evidence, while $\pi(\bm{\theta}_{i}^{k}|\bm{H})$ represents prior in the population model, controlled by the hyperparameters $\bm{H}$.
We compare the posterior distributions, $P(\bm{H}|\mathcal{D})$, based on three sets of samples: the raw samples for each event from GWOSC, the recovered samples using corresponding neural network weights and the recovered samples using corresponding analytic expressions. All three sets contain the same 10 events, and for each event the number of posterior samples equals to that of the raw samples. We use the GWPopulation package Talbot:2024yqw to perform population inference on these three sets of samples. The analyses use the two-component-primary-mass-ratio and iid-spin models, with detailed description and prior choices of the hyperparameters given in Appendix D. Comparative results for the three population-level posterior distributions are shown in Tab. 3 and Fig. 8. As shown in the figure, population inference results of raw samples and samples recovered using neural network weights match very well, validating the feasibility of our compression strategy. Result of samples recovered using analytic expressions exhibits slight deviations. This behavior may in part stem from the numerical instability of hierarchical Bayesian inference Leyde:2023iof, which can amplify small variations in the single-event posteriors, particularly for weakly constrained hyperparameters such as the mass-ratio slope.
Figure 8: Marginalized one- and two-dimensional distributions, comparing the population inference results using raw samples (blue), samples generated with our neural networks (orange) and analytic expressions (pink) of 10 GW events. Contour lines in the two-dimensional joint distributions delineate the $50\%$ and $90\%$ credible regions. The sampling is accomplished using the dynesty sampler Speagle:2019ivv and nlive is set to 2,000.
Table 3: Normalized one-dimensional Wasserstein distances for the population posteriors. The two rows show the comparisons between the population inference results using raw samples and the samples generated using neural network weights and the samples generated using analytic expressions, respectively.
Type $\alpha$ $\beta$ $_{\rm min} mm$$_{\rm max}$ $\lambda_{m}$ $\mu_{m}$ $\sigma_{m}$ $\alpha_{\chi}$ $\beta_{\chi}$ $\xi$ $\sigma$
Network Weight 0.018 0.032 0.027 0.028 0.049 0.012 0.022 0.030 0.015 0.014 0.018
Analytic Expression 0.199 0.225 0.032 0.069 0.169 0.028 0.046 0.165 0.012 0.064 0.215
@@ Acknowledgements
We thank the anonymous referee for comments. This work was supported by the Beijing Natural Science Foundation (1242018), the National Natural Science Foundation of China (123B2043, 12573042), the National SKA Program of China (2020SKA0120300), the Max Planck Partner Group Program funded by the Max Planck Society, and the High-performance Computing Platform of Peking University.
@@
Appendix A A Case Test Beyond Gaussian Families
In this appendix, we present an additional illustrative example beyond Gaussian families to more comprehensively demonstrate the effectiveness of our method in modeling diverse distributions. In Sec. III, we chose three cases parameterized by Gaussian or Gaussian mixture distributions because they provide closed-form ground truths. It is worth noting that our method remains effective beyond Gaussian families. To illustrate this, we present an additional example using an exponential-type distribution:
$P(\theta_{1},\theta_{2})=\frac{1}{\theta_{1}}e^{-\theta_{1}}e^{-\frac{\theta_{2}}{\theta_{1}}}.$
We fit this target with three mixture components per conditional and obtain the following learned expression,
[…]
$0].\displaystyle+24e^{-0.77(0.65\theta_{1}+0.63\theta_{2}+1.00)^{2}}+00$].
The reconstructed distribution is visualized in Fig. 9. This example demonstrates that, in practice, our estimator can recover the target distribution with high fidelity, serving as an additional validation of the method’s effectiveness.
Figure 9: Results of additional illustrative example. Marginalized one- and two-dimensional distributions are shown in corner plots, comparing samples resampled from fitted expressions (orange) and samples generated from the ground-truth distribution (blue). Both of them contain 40,000 samples.
Appendix B GW Events and Sampling Results
In this appendix, we give some information of binary black hole GW events in GWTC-1 (GW150914, GW151012, GW151226, GW170104, GW170608, GW170729, GW170809, GW170814, GW170818, GW170823 LIGOScientific:2018mvr), and the resampling results obtained with our compact data products.
All ten GW events used in this paper are binary black hole mergers composed of stellar-mass black holes, originally detected during the LIGO/Virgo’s first and second observing runs LIGOScientific:2018mvr. The raw posterior samples used in this paper are taken from the GWTC-2.1 catalog LIGOScientific:2021usb, which features a reanalysis of these events with state-of-the-art waveform models. This population covers a range of total masses from approximately $18.5\,M_{\odot}$ (GW170608) to $84.4\,M_{\odot}$ (GW170729). While most systems have comparable component masses, the spin characteristics show important variations. The majority of events have effective inspiral spins ($\chi_{\rm eff}$) consistent with zero. However, GW151226 stands out with a measured positive spin, providing evidence for at least one spinning black hole in that system. Furthermore, parameter uncertainties vary significantly across the events, often correlating with detection significance: the inaugural loud event GW150914 with a network SNR of $26$ has well-constrained properties, whereas the most distant event GW170729 at redshift $z\approx 0.44$ exhibits much broader posterior distributions due to its lower SNR.
The sampling results are shown in Fig. 10 and Fig.11 and the normalized one-dimensional Wasserstein distances are listed in Tab. 4. As shown in Fig. 10, the samples generated from the neural network weights exhibit close agreement with the original distribution from GWOSC. In Fig. 11, samples generated with the analytic expressions capture the overall structure of the posterior distribution, but exhibit some deviations from the original distributions. In producing the results in the figures, the data preprocessing method, neural network configurations, training processes, and training devices of all ten events are kept consistent. During resampling, the number of samples for each event is the same as that of the raw samples, and the resampling device also remains consistent.